Circular No. 5644

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER (1992y) Carolyn S. Shoemaker reports her discovery of a comet on films taken

by E. M. Shoemaker, D. H. Levy, and H. Holt with the 0.46-m Schmidt telescope at Palomar. There is a tail fanned toward the south-southwest. D. D. Balam, University of Victoria, provides a precise position obtained with the 0.25-m Schmidt telescope.

1992 UT  α_{2000}  δ_{2000}  m_{1} Observer
Oct. 25.47413 4^{h}23.^{m}5  +25^{o}45′  15.5 Shoemaker
26.35416 4 23.0  +26 04  _{”}
27.42646 4 21 16.^{s}08 +26 24 26′′.1  Balam

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence, com-

municate: “CCD photometry of the optical counterpart of the hard x-ray transient GRO J0422+32 (IAUC 5580, 5588) was obtained with the 1.2-m telescope at Haute-Provence during three different runs (Sept. 1-4, 18-21, and 30-Oct. 1) in the B, V, and I bands. Time analysis of this data set shows that a low-amplitude periodic modulation is superposed onto a gradual 0.25-mag decline in brightness, with small (< 0.1 mag) night-to-night changes in mean level. The most probable period is 4.96 hr, but the 1-day alias at 4.50-hr cannot be ruled out. The light curve exhibits two maxima and two minima per cycle and has a mean peak-to-peak amplitude of 0.04 magnitude.”

BZ URSAE MAJORIS J. A. Mattei, American Association of Variable Star Observers, reports

that this star has gone into bright outburst, as indicated by the following visual magnitude estimates: Oct. 26.29 UT, 10.5 (D. York, Abiquiu, NM); 27.25, 11.8 (York); 27.38, 12.2 (J. McKenna, Upper Montclair, NJ). This cataclysmic variable has not been this bright since it reached m_{v} = 10.3 in 1975 October, and its last confirmed outburst was at m_{v} ∼ 12 in 1991 April; Mattei encourages that high-speed photometry be undertaken.

1992 October 27 Daniel W. E. Green